# Inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities

### v1

@@ Abstract
We study nonlinear elliptic equations arising as stationary states of inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities. Working in a weighted Sobolev space that combines the homogeneous Sobolev space with a weighted Lebesgue term, we establish continuous and compact embeddings of Caffarelli–Kohn–Nirenberg type. These embeddings, together with a model that displays a natural scaling, allow us to apply the abstract critical point framework of Mercuri and Perera (2025), yielding a sequence of nonlinear eigenvalues for the associated problem. This scaling property leads to a classification of weighted power-type nonlinearities into subscaled, scaled, and superscaled regimes. Within this variational setting, we obtain broad existence and multiplicity results for equations driven by sums of weighted power nonlinearities, covering superscaled, scaled, and subscaled interactions, both in the subcritical and critical cases. We also provide a nonexistence result as a consequence of a Pohozaev-type identity. Finally, in the radial setting we employ improved radial CKN inequalities to enlarge the admissible embedding ranges. This yields strengthened radial versions of all our main results, including two-dimensional configurations with more singular weights, where no compact embeddings are available in the nonradial case.
@@ 1 Introduction
where $0<a,b<2$, $2<p<q$, and $N\geq 2$. This problem can be viewed as the search for standing wave profiles of the time-dependent INLS with two competing nonlinearities.
@@
Within this setting, we carry out a comprehensive analysis of nonlinearities given by sums of weighted power-type terms, under assumptions dictated by their asymptotic behavior relative to the scaled nonlinearity and the sequence of eigenvalues. We establish existence and multiplicity results for (1.4) across all the corresponding regimes. The precise statements of our main theorems are presented in Subsection 1.2.
@@ 1.1 Variational framework
We define the natural energy space to look for solutions to problems (1.1) and (1.4). For $N\geq 2$, $r\geq 1$ and $\eta\in(0,N)$ we denote by $L^{r}_{\eta}(\mathbb{R}^{N})$ the weighted Lebesgue space $L^{r}(\mathbb{R}^{N};|x|^{-\eta}dx)$. Recall that $L^{r}_{\eta}(\mathbb{R}^{N})$ is a Banach space with the norm
@@
and it is uniformly convex if $r>1$. Considering $b\in(0,N)$ and $q\geq 1$, we define
@@
As usual, a key step in formulating the energy functional associated with our problem is to determine into which Lebesgue spaces the underlying Sobolev-type space embeds. For $N\geq 2$, $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<N$ let us define
@@
Note that $2^{*}_{b,q,\eta}\leq\min\{q,2^{*}_{\eta}\}<2^{*}$ when $b\leq\eta$, while for $0\leq\eta<b$ and $N\geq 3$ one has $q<2^{*}_{b,q,\eta}<2^{*}_{\eta}\leq 2^{*}$.
@@ 1.2 Main theorems
Notice that if $(\eta,r)$ is subscaled (scaled, superscaled) then
@@
Thus, the existence of a subscaled (scaled) pair $(\eta,r)$ requires $\eta>b$ ($\eta\geq b$), since for $N\geq 3$ one has
$q+\frac{b-\eta}{\delta}<\frac{q(N-\eta)}{N-b}<2^{*}_{b,q,\eta},\quad\text{for all}\quad\eta\in[0,b).$
@@
We begin by proving existence of nontrivial solutions for subcritical superscaled problems. Our first theorem reads
@@
for some constants $c,c_{1},c_{2}>0$, and superscaled pairs $(\eta,r_{1})$, $(\eta,r_{2})$ and $(\eta,r)$, with $2^{*}_{b,q,\eta}<r_{1},r_{2}<2^{*}_{\eta}$ and $q<r<2^{*}_{\eta}$. Then, problem (1.4) has a nontrivial solution. Moreover, if $\lambda$ is not an eigenvalue of (1.1), this solution has positive energy.
@@
$\eta_{2}\in[0,\eta_{1})\quad\text{and}\quad 2^{*}_{\eta_{1}}\leq r_{2}<2^{*}_{\eta_{2}}.$ (1.11)
$\eta_{2}\in[\eta_{1},2)\,\,(\eta_{2}=\eta_{1}\,\,\text{only if}\,\,\eta_{1}>b)\quad\text{and}\quad r_{2}<2^{*}_{\eta_{2}},\quad\text{with}\,\,r_{2}>q+\frac{(b-\eta_{2})(r_{1}-q)}{b-\eta_{1}}\quad\text{if}\quad\eta_{2}<b.$ (1.12)
@@
Assume $N\geq 3$ and ($\mathcal{H}$). Consider $f$ as in (1.9). Then, for every $\lambda\in\mathbb{R}$, $\eta_{1}$ satisfying (1.10) and $(\eta_{2},r_{2})$ superscaled satisfying one of the conditions (1.11)-(1.12), and $m\in\mathbb{N}$ there exists $\mu_{m}>0$ such that, for all $\mu>\mu_{m}$, problem (1.4) admits at least $m$ distinct pairs of nontrivial solutions with positive energy. In particular, the number of these solutions tends to infinity as $\mu\to\infty$.
@@
Assume ($\mathcal{H}$). For a subscaled pair $(\eta,r)$, with $b<\eta<2$, consider
@@
For the next Theorem, recall the notation given in (1.7).
@@
$\ell(\eta,r)<\ell(\eta_{1},r_{1}),\quad\mbox{for}\,\,\lambda\leq 0\quad\mbox{or}\quad(\eta_{1},r_{1})\text{ superscaled }\quad{for}\,\,\lambda>0.$
@@
where $(\eta,r)$ is subscaled and $0\leq\eta_{1}<2$. The there exists a $\mu^{*}>0$ such that problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels with $\Phi(u_{k})\nearrow 0$ for each $\mu\in(0,\mu^{*})$.
@@
The structure of the remainder of the paper is as follows: In Section 2 we give some preliminary results such as the embedding theorems that justify Definition 1, and a Pohozaev Identity. Section 3 is devoted to the abstract framework of Mercuri and Perera [23]. We verify that our framework satisfy the structural hypotheses of the theory, and we construct the nonlinear eigenvalues $(\lambda_{k})$ of problem (1.1) using the Fadell–Rabinowitz cohomological index. Section 4 contains all proofs of the main results. We first establish general Palais–Smale and boundedness criteria. These preliminary tools are then used to treat, in a unified manner, some classes of nonlinearities considered in the paper: superscaled perturbations in the subcritical and critical cases; combinations of scaled, subscaled, and superscaled terms; and problems involving the nonlinear eigenvalues $(\lambda_{k})$. Section 5 is devoted to the radial setting. Using the improved radial Caffarelli–Kohn–Nirenberg inequalities, we obtain strictly larger embedding ranges for $E_{b,\mathrm{rad}}^{q}$ and extend the admissible regimes for $(\eta,r)$, particularly in the case $\eta<b$ and in dimension $N=2$, where no nonradial compact embeddings are available. We explain how all main existence and multiplicity theorems admit strengthened radial versions under these improved embedding intervals. Finally, Section 6 contains an appendix where we prove the density of smooth functions in the underlying functional space used throughout this paper.
@@ 2.1 Embeddings in weighted Lebesgue spaces
Due to the Caffarelli - Kohn - Nirenberg inequality (see [7]), for fixed numbers
@@
[22, Proposition 24] Let $N\geq 1$, $q\geq 1$, $r\geq 1$, $0<\theta<1$, $0<b<N$ and $\gamma\in\mathbb{R}$. Define $\sigma$ by (2.1) and assume that $\sigma<0$, (2.3) holds and
@@
Suppose $N\geq 2$, let $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<N$. For $N=2$, assume, additionally, that $b\leq\eta$. Fix $r>0$ such that
@@
Moreover, if we assume $\eta\leq b$ we can consider $r=2^{*}_{b,q,\eta}$ and, if $\eta\leq 2$ and $N\geq 3$, the other endpoint $r=2^{*}_{\eta}$. Then, it holds
@@
In particular, we have $E^{q}_{b}\hookrightarrow L_{\eta}^{r}(\mathbb{R}^{N})$ for every $r\geq 1$ in the above interval. Moreover, for $N=2$ with $\eta>b$, or for $N\geq 3$, these embeddings are compact except at the endpoints of the interval.
@@
Fix $r$ in the closure of one of the intervals and choose $\theta\in[0,1]$ so that (2.5) holds for this value of $r$. Consider $\gamma=-\eta/r$ so that (2.3) can be written as (2.5). Let $\sigma\in\mathbb{R}$ be given such that $\gamma=\theta\sigma-(1-\theta)b/q$, as in (2.1). Then, if $\theta=0$ we see that this expression of $\gamma$ and (2.5) imply
$\frac{\eta}{r}=-\gamma=\frac{b}{q}\quad\text{and}\quad\frac{1}{r}=\frac{N-b}{q(N-\eta)}\qquad\Longrightarrow\qquad\eta=b\,\,\,\text{and}\,\,\,r=q.$
If $\theta>0$, due to (2.5) we get
@@
where $\theta\in(0,1)$ if $\eta>b$ and $\theta\in[0,1)$ if $\eta=b$. On the other hand, for $N\geq 3$ we get $\sigma\leq 0$ for any $\theta\in(0,1]$ if $b\leq\eta$ while for $0\leq\eta<b$ it holds
$\sigma\leq 0\quad\Longleftrightarrow\quad\theta\geq\theta_{1}:=\frac{2N(b-\eta)}{\eta q(N-2)+2N(b-\eta)}\in(0,1].$
By placing $\theta=\theta_{1}$ in (2.5) we obtain $r=2^{*}_{b,q,\eta}$. Moreover, when $\theta=1$ and $\eta\leq 2$ we get $\sigma\geq-1$. So, $\sigma$ satisfies (2.4) and then (2.2) occurs for $r$ between $2^{*}_{b,q,\eta}$ and ${2(N-\eta)}/{(N-2)}$, including ${2(N-\eta)}/{(N-2)}$ only in case $\eta\leq 2$, and including $2^{*}_{b,q,\eta}$} only in case $\eta\leq b$.
Under these conditions, we conclude that (2.6) follows by density.
It remains to prove that these embeddings are compact (except at the endpoints). Let $(u_{n})$ be a bounded sequence of functions in $E^{q}_{b}$ with $u_{n}\in C_{c}^{\infty}(\mathbb{R}^{N})$. This means that the sequences $(\|\nabla u_{n}\|_{2})$ and $\bigl(\|\,|x|^{-b/q}u_{n}\|_{q}\bigr)$ are bounded. Then, if $N=2$ with $\eta>b$ or $N\geq 3$, fixing $r\geq 1$ in one of those (open) intervals and setting $\gamma=-\eta/r$, all the chosen parameters satisfy the assumptions of Proposition 2.1. Notice that the space $W^{1,2}(B_{1})\cap L^{q}(B_{1})$ is compactly embedded into $L^{k}(B_{1})$ for every $1\leq k<\max\{q,\,2N/(N-2)\}$ (or $k\geq 1$ if $N=2$), and $r$ lies within this range. Hence, $\bigl(|x|^{-\eta/r}u_{n}\bigr)$ admits a convergent subsequence in $L^{r}(\mathbb{R}^{N})$, which completes the proof. ∎
@@
For $N\geq 2$ we verify that
$2^{*}_{b,q,a}<p\quad\Longleftrightarrow\quad q[(N-2)(q-2)-2(2-b)]<p[(N-2)(q-2)-2(2-b)]$
and for $N\geq 3$ we get
$p<2^{*}_{a}\quad\Longleftrightarrow\quad p[(N-2)(q-2)-2(2-b)]<2[(N-2)(q-2)-2(2-b)].$
Thus, since $p\in(2,q)$ and $q<2^{*}_{b}$, we obtain $2^{*}_{b,q,a}<p<2^{*}_{a}.$

### v2

@@ Abstract
We study nonlinear elliptic equations that arise as stationary states of inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities. The model involves the Laplacian combined with weighted power-type terms and naturally leads to a variational formulation in a weighted Sobolev space obtained from the intersection of the homogeneous Sobolev space with a weighted Lebesgue space. Using sharp weighted Sobolev and Caffarelli–Kohn–Nirenberg type inequalities, we establish continuous and compact embeddings of this space into suitable weighted Lebesgue spaces. These embedding results, together with a natural scaling structure of the model, allow us to apply the abstract critical point framework of Mercuri and Perera (2026), yielding a sequence of nonlinear eigenvalues for the associated problem via a min–max scheme based on the Fadell–Rabinowitz cohomological index. Within this framework we obtain a broad collection of existence and multiplicity results for equations driven by sums of weighted power nonlinearities, covering interactions in both subcritical and critical cases. We also establish a nonexistence result derived from a Pohozaev-type identity. Finally, we analyze the radial setting, where improved radial Caffarelli–Kohn–Nirenberg inequalities allow us to enlarge some of the admissible embedding ranges. This leads to strengthened radial versions of our main results.
@@ 1 Introduction
where $0\leq b<a<2$, $2<p<q$, and $N\geq 2$. This problem can be viewed as the search for standing wave profiles of the time-dependent INLS with two competing nonlinearities.
@@
Within this setting, we carry out a comprehensive analysis of nonlinearities given by sums of weighted power-type terms, under assumptions dictated by their asymptotic behavior relative to the scaled nonlinearity and the sequence of eigenvalues. We establish existence and multiplicity results for (1.4) across all the corresponding regimes. The precise statements of our main theorems are presented in Subsection 1.2. It is worth mentioning that some of these results are also new in the case $b=0$, where the singular weights appear only in the nonlinearities and not in the operator itself. This situation is particularly interesting in the radial setting, where our approach provides new insights into such problems. See Section 5.
@@ 1.1 Variational framework
We define the natural energy space to look for solutions to problems (1.1) and (1.4). For $N\geq 2$, $r\geq 1$ and $\eta\in[0,N)$ we denote by $L^{r}_{\eta}(\mathbb{R}^{N})$ the weighted Lebesgue space $L^{r}(\mathbb{R}^{N};|x|^{-\eta}dx)$. Clearly, $L^{r}_{0}(\mathbb{R}^{N})=L^{r}(\mathbb{R}^{N})$. Recall that $L^{r}_{\eta}(\mathbb{R}^{N})$ is a Banach space with the norm
@@
and it is uniformly convex if $r>1$. Considering $b\in[0,N)$ and $q\geq 1$, we define
@@
As usual, a key step in formulating the energy functional associated with our problem is to determine into which Lebesgue spaces the underlying Sobolev-type space embeds. For $N\geq 2$, $0\leq b<2<q<2^{*}_{b}$ and $0\leq\eta<N$ let us define
@@
Note that $2^{*}_{b,q,\eta}\leq\min\{q,2^{*}_{\eta}\}<2^{*}$ when $b\leq\eta$, while for $N\geq 3$ and $\eta\in(0,b)$ one has $q<2^{*}_{b,q,\eta}<2^{*}_{\eta}<2^{*}$ and for $\eta=0<b$ it holds $q<2^{*}_{b,q,0}=2^{*}_{0}=2^{*}$.
@@ 1.2 Main theorems
Notice that $(\eta,r)$ is subscaled (scaled, superscaled) if and only if
@@
Thus, the existence of a subscaled (scaled) pair $(\eta,r)$ requires $\eta>b$ ($\eta\geq b$), since one has
$q+\frac{b-\eta}{\delta}<\frac{q(N-\eta)}{N-b}\ \left(<2^{*}_{b,q,\eta}\text{ for }N\geq 3\right),\quad\text{for all}\quad\eta\in[0,b).$
@@
We are now ready to state the main results of this paper. We begin by considering subcritical superscaled problems and establishing the existence of nontrivial solutions. Our first theorem reads as follows.
@@
for some constants $c,c_{1},c_{2}>0$, and superscaled pairs $(\eta,r_{1})$, $(\eta,r_{2})$ and $(\eta,r)$, with $2^{*}_{b,q,\eta}<r_{1},r_{2}<2^{*}_{\eta}$ and $q<r<2^{*}_{\eta}$. Then, problem (1.4) has a nontrivial solution. If $\lambda$ is not an eigenvalue of (1.1), then (1.4) has a solution with positive energy.
@@
$\eta_{2}\in(0,\eta_{1})\quad\text{and}\quad 2^{*}_{\eta_{1}}\leq r_{2}<2^{*}_{\eta_{2}}.$ (1.11)
$\eta_{2}\in[\eta_{1},2)\,\,(\eta_{2}=\eta_{1}\,\,\text{only if}\,\,\eta_{1}>b)\quad\text{and}\quad r_{2}<2^{*}_{\eta_{2}},\quad\text{with}\,\,r_{2}>q+(2^{*}_{\eta_{1}}-q)\frac{(b-\eta_{2})}{b-\eta_{1}}\quad\text{if}\quad\eta_{2}<b.$ (1.12)
@@
Assume $N\geq 3$ and ($\mathcal{H}$). Consider $f$ as in (1.9). Then, for every $\lambda\in\mathbb{R}$, $\eta_{1}$ satisfying (1.10), $(\eta_{2},r_{2})$ superscaled satisfying one of the conditions (1.11)-(1.12), and $m\in\mathbb{N}$, there exists $\mu_{m}>0$ such that, for all $\mu>\mu_{m}$, problem (1.4) admits at least $m$ distinct pairs of nontrivial solutions with positive energy. In particular, the number of these solutions tends to infinity as $\mu\to\infty$.
@@
Assume ($\mathcal{H}$). For a subscaled pair $(\eta,r)$, consider
@@
For the next theorem, recall the notation $\ell(\eta,r)=r\delta+\eta-N$ and $\ell=\ell(b,q)=\ell(a,p)$, given in (1.7).
@@
$\ell(\eta,r)<\ell(\eta_{1},r_{1})\quad\mbox{for}\,\,\lambda\leq 0\quad\mbox{or}\quad(\eta_{1},r_{1})\text{ superscaled }\quad{for}\,\,\lambda>0.$
@@
An interesting feature of the result above is that it also covers the case $b=\eta_{1}=0$, despite the fact that the embeddings $E^{q}_{0}\hookrightarrow L^{r_{1}}(\mathbb{R}^{N})$ are not compact for any $q\leq r_{1}\leq 2^{*}$ (or $q\leq r_{1}<\infty$ if $N=2$).
@@
where $(\eta,r)$ is subscaled and $0\leq\eta_{1}<2$. Then there exists $\mu^{*}>0$ such that problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels with $\Phi(u_{k})\nearrow 0$ for each $\mu\in(0,\mu^{*})$.
@@
The structure of the remainder of the paper is as follows: In Section 2 we give some preliminary results such as the embedding theorems that justify Definition 1, and a Pohozaev Identity. Section 3 is devoted to the abstract framework of Mercuri and Perera [23]. We verify that our framework satisfy the structural hypotheses of the theory, and we construct the nonlinear eigenvalues $(\lambda_{k})$ of problem (1.1) using the Fadell–Rabinowitz cohomological index. Section 4 contains all proofs of the main results. We first establish general Palais–Smale and boundedness criteria. These preliminary tools are then used to treat, in a unified manner, some classes of nonlinearities considered in the paper: superscaled perturbations in the subcritical and critical cases; combinations of scaled, subscaled, and superscaled terms; and problems involving the nonlinear eigenvalues $(\lambda_{k})$. Section 5 is devoted to the radial setting. Using the improved radial Caffarelli–Kohn–Nirenberg inequalities, we obtain strictly larger embedding ranges for the radial space $E^{q}_{b,\mathrm{rad}}$ by extending the admissible regimes for $(\eta,r)$. When $N\geq 3$, the range of admissible exponents can be enlarged in the case $0\leq\eta<b$. In dimension $N=2$, where the nonradial theory does not allow $\eta<b$, the radial embeddings yield a new interval of admissible exponents for $0\leq\eta<b$. Moreover, in any dimension we may treat the case $\eta=b=0$, obtaining subcritical compact embeddings that are not available in the nonradial setting. We then show how all main existence and multiplicity results admit strengthened radial versions under these improved embedding intervals. Finally, Section 6 contains an appendix where we prove the density of smooth functions in the underlying functional space used throughout this paper.
@@ 2.1 Embeddings in weighted Lebesgue spaces
Consider $b\in[0,N)$. Due to the Caffarelli - Kohn - Nirenberg inequality (see [7]), for fixed numbers
@@
[22, Proposition 24] Let $N\geq 1$, $q\geq 1$, $r\geq 1$, $0<\theta<1$, $0\leq b<N$ and $\gamma\in\mathbb{R}$. Define $\sigma$ by (2.1) and assume that $\sigma<0$, (2.3) holds and
@@
Suppose $N\geq 2$, let $0\leq b<2<q<2^{*}_{b}$ and $0\leq\eta<N$. For $N=2$, assume, additionally, that $b\leq\eta$. Fix $r>0$ such that
@@
In addition, if we assume $\eta\leq b$ we can consider $r=2^{*}_{b,q,\eta}$ and, if $\eta\leq 2$ and $N\geq 3$, the other endpoint $r=2^{*}_{\eta}$. Then, it holds
@@
In particular, we have $E^{q}_{b}\hookrightarrow L_{\eta}^{r}(\mathbb{R}^{N})$ for every $r\geq 1$ in the above interval. Moreover, for $N\geq 3$ with $\eta>0$, or $N=2$ with $\eta>b$, these embeddings are compact except at the endpoints of the interval.
@@
Fix $r\in[2^{*}_{b,q,\eta},2^{*}_{\eta}]$ and choose $\theta\in[0,1]$ so that (2.5) holds for this value of $r$. Consider $\gamma=-\eta/r$ so that (2.3) can be written as (2.5). Let $\sigma\in\mathbb{R}$ be given such that $\gamma=\theta\sigma-(1-\theta)b/q$, as in (2.1). Then, if $\theta=0$ we see that this expression of $\gamma$ and (2.5) imply $\eta=b$ and $r=q$. If $\theta>0$, due to (2.5) we get
@@
where $\theta\in(0,1)$ if $\eta>b$ or $\theta\in[0,1)$ if $\eta=b$. On the other hand, for $N\geq 3$ we get $\sigma\leq 0$ for any $\theta\in(0,1]$ if $b\leq\eta$, while for $0\leq\eta<b$ it holds $\sigma\leq 0$ if, and only if, $\theta$ satisfies
$\theta\geq\theta_{1}:=\frac{2N(b-\eta)}{\eta q(N-2)+2N(b-\eta)}\in(0,1].$
Notice that, in this case $0\leq\eta<b$, $\theta\geq\theta_{1}$ is equivalent to $r\geq 2^{*}_{b,q,\eta}$. Moreover, when $\theta=1$ and $\eta\leq 2$ we get $\sigma\geq-1$. So, $\sigma$ satisfies (2.4) and then (2.2) occurs for $r$ between $2^{*}_{b,q,\eta}$ and $2^{*}_{\eta}$, including $2^{*}_{\eta}$ only in case $\eta\leq 2$, and including $2^{*}_{b,q,\eta}$} only in case $\eta\leq b$. Under these conditions, we conclude that (2.6) follows by density.
It remains to prove that these embeddings are compact (except at the endpoints). Let $(u_{n})$ be a bounded sequence of functions in $E^{q}_{b}$ with $u_{n}\in C_{c}^{\infty}(\mathbb{R}^{N})$. This means that the sequences $(\|\nabla u_{n}\|_{2})$ and $\bigl(\|\,|x|^{-b/q}u_{n}\|_{q}\bigr)$ are bounded. Then, if $N\geq 3$ and $\eta>0$, or $N=2$ with $\eta>b$, fixing $r\geq 1$ in $(2^{*}_{b,q,\eta},2^{*}_{\eta})$ and setting $\gamma=-\eta/r$, all the chosen parameters satisfy the assumptions of Proposition 2.1. Notice that the space $W^{1,2}(B_{1})\cap L^{q}(B_{1})$ is compactly embedded into $L^{k}(B_{1})$ for every $1\leq k<2^{*}$ (or $k\geq 1$ if $N=2$), and $r$ lies within this range. Hence, $\bigl(|x|^{-\eta/r}u_{n}\bigr)$ admits a convergent subsequence in $L^{r}(\mathbb{R}^{N})$, which completes the proof. ∎
@@
For $N\geq 2$ we verify that $2^{*}_{b,q,a}<p$ if, and only if,
q[q(N2)2(Nb)]<p[q(N2)2(Nb)].q[q(N-2)-2(N-b)]<p[q(N-2)-2(N-b)].
Moreover, for $N\geq 3$ we get $p<2^{*}_{a}$ if, and only if,
p[q(N2)2(Nb)]<2[q(N2)2(Nb)].p[q(N-2)-2(N-b)]<2[q(N-2)-2(N-b)].
Thus, since $p\in(2,q)$ and $q<2^{*}_{b}$, we obtain $2^{*}_{b,q,a}<p<2^{*}_{a}.$ The result then follows from Lemma 2.2. ∎
Another useful compact embedding, which will be used throughout the paper, is given in the next corollary. Its proof follows directly from the definition of an admissible subscaled pair $(\eta,r)$ (Definition 1) together with Lemma 2.2.
Corollary 2.4.
Assume ($\mathcal{H}$). If $(\eta,r)$ is a subscaled pair, then the embedding $E^{q}_{b}(\mathbb{R}^{N})\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ is compact.